Angles 1 and 2 are complimentary. If angle 1 = 5x + 15 and angle 2 = 10x. The measure of angle 1 is:
Answer by nyc_function(2741)

Angles 1 and 2 are complementary.


If this is true, then angle 1 + angle 2 = 90.


5x + 15 + 10x = 90


15x + 15 = 90


15x = 90 - 15


15x = 75


x = 75/15


x = 5


angle 1 = 5x + 15


Let x = 5 and solve for angle 1.


angle 1 = 5(5) + 15


angle 1 = 25 + 15


angle 1 = 40 degrees.